# UniAP: Unifying Inter- and Intra-Layer Automatic Parallelism by Mixed Integer Quadratic Programming

### v4

@@ Appendix C QIP formulation for intra-layer-only parallelism
$\sum_{u\in{\mathbb{V}}}{\boldsymbol{\mathit{S}}}_{u}^{\mathsf{T}}{\boldsymbol{\mathit{A}}}_{u}+\sum_{\langle u,v\rangle\in{\mathbb{E}}}{\boldsymbol{\mathit{S}}}_{u}^{\mathsf{T}}{\boldsymbol{\mathit{R}}}_{uv}{\boldsymbol{\mathit{S}}}_{v}=p_{1}.~{}$ (11)
@@
$\sum_{u\in V}{\boldsymbol{\mathit{S}}}_{u}^{\mathsf{T}}{\boldsymbol{\mathit{M}}}_{u}\leqslant m.~{}$ (12)
@@
Appendix D Visualization for the candidate solution
@@
In this section, we proceed to visually represent a potential solution for UOP. Given a deep learning model $\mathcal{G}$, pipeline parallel size $pp\_size$, and number of micro-batches $c$, UniAP will determine layer placement ${\boldsymbol{\mathit{P}}}$ for inter-layer parallelism and parallel strategy ${\boldsymbol{\mathit{S}}}$ for intra-layer parallelism using an off-the-shelf solver. As Figure 6 shows, the solver is optimizing a three-layer model with two pipeline stages, each assigned four GPUs. At this time, a candidate solution could be
${\boldsymbol{\mathit{P}}}=\begin{bmatrix}1&0\\ 1&0\\ 0&1\end{bmatrix},~{}{\boldsymbol{\mathit{S}}}=\begin{bmatrix}1&0&0\\ 0&1&1\\ 0&0&0\\ \vdots&\vdots&\vdots\\ 0&0&0\end{bmatrix}.$ (13)
Here, the $u$-th row of matrix ${\boldsymbol{\mathit{P}}}$ denotes the placement strategy for layer $u$, where ${\boldsymbol{\mathit{P}}}_{ui}=1$ signifies the placement of layer $u$ on stage $i$, while 00 indicates otherwise. For example, ${\boldsymbol{\mathit{P}}}_{l_{0}}=\left[1,~{}0\right]$ denotes the placement of layer $l_{0}$ on pipeline stage 1. Additionally, the $u$-th column of matrix ${\boldsymbol{\mathit{S}}}$ denotes the selected intra-layer parallel strategy for layer $u$, where ${\boldsymbol{\mathit{S}}}_{uj}=1$ denotes the selection of the $j$-th strategy from the intra-layer parallel strategy set. For example, ${\boldsymbol{\mathit{S}}}_{l_{0}}=\left[1,~{}0,~{}0,~{}\cdots,~{}0\right]^{\mathsf{T}}$ indicates that layer $l_{0}$ will adopt only the DP strategy, while ${\boldsymbol{\mathit{S}}}_{l_{1}}=\left[0,~{}1,~{}0,~{}\cdots,~{}0\right]^{\mathsf{T}}$ indicates that layer $l_{1}$ will employ a strategy where DP is performed on GPU 0, 1 and GPU 2, 3, and TP is performed across these two GPU groups.
There exist numerous combinations of ${\boldsymbol{\mathit{P}}}$ and ${\boldsymbol{\mathit{S}}}$. The off-the-shelf solver will automatically search for the optimal solution given pipeline parallel size $pp\_size$ and the number of micro-batches $c$. By solving the MIQP expression and enumerating every possible $pp\_size$ and $c$ in the UOP process, UniAP will ultimately derive an optimal parallel strategy for the deep learning model within the current hardware environment.
Appendix E Experiment detail
@@ Appendix F Case study: BERT-Huge
Appendix G Scaling to larger clusters and larger models
We adapt UniAP to a cloud cluster to further scale UniAP to larger clusters and larger models. The experiments are conducted on 8 nodes. Each of them has 1 Intel Xeon Phi Processor 7285, 4 16GB PCIe DCU, 128GB DDR4 DRAM, and 200Gb Infiniband interconnection network. We name this environment as EnvE. It is worth noting that DCU is a kind of GPU, which is different from NVIDIA GPUs. We select MP methods including Megatron [18] and DeepSpeed (ZeRO-3) [17] as our baseline methods.
In the experiment, we evaluate these baselines and UniAP with Llama-7B and Llama-13B [6]. We set the mini-batch size as 8 for Llama-7B and 4 for Llama-13B. The remaining configurations align with those for Llama-7B in Section 4, such as activating FP16 mixed precision training.
We focus on two metrics: training throughput and strategy optimization time. For UniAP, the former metric is calculated by averaging throughput from the 10th to the 60th iteration of training, while the later metric is determined by measuring the time of the UOP (Section 3.4). Please note that Megatron [18] and DeepSpeed [17] do not automatically optimize parallel strategies. Consequently, we undertake an exhaustive approach, examining every combination of parallel strategies and the corresponding parallel configurations they accommodate. For each combination, we assess the average throughput over the 10th to 60th training iteration, ultimately adopting the fastest strategy as their optimization result. The total time spent in the process is recorded as their strategy optimization time.
Table 4: Training throughput and strategy optimization time on EnvE.
Megatron DeepSpeed UniAP Megatron DeepSpeed UniAP
  • 1)
    SOL×$\times$: No solution after strategy optimization.
As shown in Table 4, UniAP consistently identifies the fastest parallel strategies while expending considerably less time on strategy optimization compared to Megatron (about 157$\times$ faster for Llama-7B and 77$\times$ faster for Llama-13B). Please note that the strategy optimization time Megatron needs for Llama-7B surpasses that for Llama-13B. We attribute this phenomenon to the variations in the mini-batch size. Specifically, the escalation from a mini-batch size of 4 (employed in Llama-13B) to 8 (employed in Llama-7B) leads to a rise in the number of candidate parallel strategies, as well as the parallel strategies that will successfully train the model. As a result, the strategy optimization time will become increasingly long when the mini-batch size exceeds 8 in the pretraining scenario [2, 6, 7]. In this situation, UniAP will identify the optimal parallel strategy much faster.
Top-11) Top-22) Slowest3) Median4)
Llama-7B 8 2.01 1.92 0.22 0.82 41 64
Llama-13B 4 0.82 0.58 0.27 0.42 42 48
  • 1)
    Top-1: The training throughput achieved by the fastest parallel strategy.
  • 2)
    Top-2: The training throughput achieved by the second fastest parallel strategy.
  • 3)
  • 4)
  • 5)
  • 6)
    #candidate: The total number of candidate parallel strategies.
We further examine the statistics on the candidate parallel strategies for Megatron, as shown in Table 5. When the users are unable to identify the fastest parallel strategy out of hundreds of candidates, they may be required to sacrifice at least 4.4% of the training throughput for Llama-7B and 29.2% for Llama-13B. Even if they successfully eliminate all infeasible parallel strategies, which constitute the majority of candidates (Table 5), they still only have a 50% chance of identifying a parallel strategy with training throughput better than 0.82 samples/s for Llama-7B and 0.42 samples/s for Llama-13B. Consequently, it is challenging for users to identify the fastest parallel strategy for larger clusters and larger models. In such circumstances, UniAP is able to demonstrate its capacity to scale and accommodate both larger clusters and larger models.

### v5

@@ Appendix B QIP formulation for intra-layer-only parallelism
$\sum_{u\in\mathbb{V}}\textbf{{S}}_{u}^{\mathsf{T}}\textbf{{A}}_{u}+\sum_{\langle u,v\rangle\in\mathbb{E}}\textbf{{S}}_{u}^{\mathsf{T}}\textbf{{R}}_{uv}\textbf{{S}}_{v}=p_{1}.~{}$ (10)
@@
$\sum_{u\in V}\textbf{{S}}_{u}^{\mathsf{T}}\textbf{{M}}_{u}\leqslant m.~{}$ (11)
@@
Appendix C Visualization for the candidate solution
@@
In this section, we proceed to visually represent a potential solution for UOP. Given a deep learning model $\mathcal{G}$, pipeline parallel size $pp\_size$, and number of micro-batches $c$, UniAP will determine layer placement P for inter-layer parallelism and parallel strategy S for intra-layer parallelism using an off-the-shelf solver. As Figure 7 shows, the solver is optimizing a three-layer model with two pipeline stages, each assigned four GPUs. At this time, a candidate solution could be
$\textbf{{P}}=\begin{bmatrix}1&0\\ 1&0\\ 0&1\end{bmatrix},~{}\textbf{{S}}=\begin{bmatrix}1&0&0\\ 0&1&1\\ 0&0&0\\ \vdots&\vdots&\vdots\\ 0&0&0\end{bmatrix}.$ (12)
Here, the $u$-th row of matrix P denotes the placement strategy for layer $u$, where $\textbf{{P}}_{ui}=1$ signifies the placement of layer $u$ on stage $i$, while 00 indicates otherwise. For example, $\textbf{{P}}_{l_{0}}=\left[1,~{}0\right]$ denotes the placement of layer $l_{0}$ on pipeline stage 1. Additionally, the $u$-th column of matrix S denotes the selected intra-layer parallel strategy for layer $u$, where $\textbf{{S}}_{uj}=1$ denotes the selection of the $j$-th strategy from the intra-layer parallel strategy set. For example, $\textbf{{S}}_{l_{0}}=\left[1,~{}0,~{}0,~{}\cdots,~{}0\right]^{\mathsf{T}}$ indicates that layer $l_{0}$ will adopt only the DP strategy, while $\textbf{{S}}_{l_{1}}=\left[0,~{}1,~{}0,~{}\cdots,~{}0\right]^{\mathsf{T}}$ indicates that layer $l_{1}$ will employ a strategy where DP is performed on GPU 0, 1 and GPU 2, 3, and TP is performed across these two GPU groups.
There exist numerous combinations of P and S. The off-the-shelf solver will automatically search for the optimal solution given pipeline parallel size $pp\_size$ and the number of micro-batches $c$. By solving the MIQP expression and enumerating every possible $pp\_size$ and $c$ in the UOP process, UniAP will ultimately derive an optimal parallel strategy for the deep learning model within the current hardware environment.
Appendix D Experiment detail
@@
Appendix E Estimation accuracy
Some variables in UniAP and other AP methods are estimated values rather than actual running values. The TPI (inverse of training throughput) returned by UniAP and other AP methods is one of them. Accurate estimation for TPI or training throughput is crucial for evaluating candidate parallel strategies and ensuring the optimality of the solution. To quantify the accuracy of the estimated training throughput, we introduce a metric called relative estimation error (REE) $e$ for training throughput:
$e(T,\hat{T})=\frac{|T-\hat{T}|}{T}\times 100\%,$ (13)
where $T$ is the actual training throughput and $\hat{T}$ is the estimated training throughput.
Figure 8: Relative estimation error.
We evaluate the optimal parallel strategies obtained from EnvA and EnvB and visualize the REE of UniAP in Figure 8. The results show that UniAP achieves an average REE of 3.59%, which is relatively small. In contrast, the average REE for Galvatron [25] in our experiments is 11.17%, which is larger than that of UniAP.